The focal length of the objective of a telescope is 75.0 cm. While the eyepiece has a focal length of 6.00 cm, what is the magnifying power of this telescope?

The answer is 12.5 I just need the work.
Answer #1

Focal length of the objective of a telescope, fo= 75.0cm

Focal length of the eyepiece of a telescope, fe= 6.0cm

The angular magnification of the telescope is given by

M= - \frac{f_o}{f_e}

using given values in above,

  M= - \frac{75.0}{6.0}= -12.5

Then magnitude of magnification of telescope will be

|M|= 12.5(ANS)